OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (6, -9) and has slope -7/6 y=mx+b y=-7/6x+b -9=-7/6(6)+b -9=42/6+b -9=7+b -2=b Did i do this coreectly

OpenStudy (anonymous):

-9=-42/6+b -9=-7+b b=-2 thats correct

OpenStudy (anonymous):

when u wrote it up there u forgot the negative sign on the 42/6 and the 7 in "-9=7+b" but u got the right answer so i assume u did it the right way when u wrote it on ur paper or something

OpenStudy (anonymous):

y=(-7/6)x-2

OpenStudy (anonymous):

ok thank you